A banana has a mean weight of 120 grams with a standard deviation of 3 grams. How many standard deviations below the mean is a banana that is 111 grams?
1
Expert's answer
2020-11-01T17:37:54-0500

"Z=\frac{X-\mu}{\sigma}"

"=\frac{111-120}{3}=-3"

A banana that is 111 is 3 standard deviations below the mean.